# SANSA: Unleashing the Hidden Semantics in SAM2 for Few-Shot Segmentation

### v1

@@ Abstract
Few-shot segmentation aims to segment unseen object categories from just a handful of annotated examples. This requires mechanisms that can both identify semantically related objects across images and accurately produce segmentation masks. We note that Segment Anything 2 (SAM2), with its prompt-and-propagate mechanism, offers both strong segmentation capabilities and a built-in feature matching process. However, we show that its representations are entangled with task-specific cues optimized for object tracking, which impairs its use for tasks requiring higher level semantic understanding. Our key insight is that, despite its class-agnostic pretraining, SAM2 already encodes rich semantic structure in its features. We propose SANSA (Semantically AligNed SegmentAnything 2), a framework that makes this latent structure explicit, and repurposes SAM2 for few-shot segmentation through minimal task-specific modifications. SANSA achieves state-of-the-art performance on few-shot segmentation benchmarks specifically designed to assess generalization, outperforms generalist methods in the popular in-context setting, supports various prompts flexible interaction via points, boxes, or scribbles, and remains significantly faster and more compact than prior approaches. Code is available at https://github.com/ClaudiaCuttano/SANSA.
@@ 1 Introduction
Segmenting images is a core problem in Computer Vision, yet achieving high-quality results typically requires extensive human effort to annotate pixel-level masks. Moreover, conventional semantic segmentation methods  [10, 20, 29, 12] struggle to generalize to unseen categories, limiting their applicability in open-world scenarios. Inspired by the human ability to recognize novel objects from just a few examples, few-shot segmentation (FSS) [33, 62, 36, 53] has emerged as a paradigm that leverages a small set of labeled reference samples to guide the segmentation of target images containing arbitrary, previously unseen classes.
To this end, recent work has turned to Visual Foundation Models (VFMs) [7], which offer rich visual representations and strong generalization capabilities [46, 72]. A natural approach [37, 73, 56] is to decouple the FSS task into two stages: feature matching followed by promptable segmentation. This is typically achieved by combining DINOv2 [45], known for its strong semantic correspondence capabilities [76, 39, 75, 57], with Segment Anything Model [30], which excels at producing high-quality segmentation masks [30, 79, 77]. While effective, these modular approaches add computational overhead and require prompt engineering to coordinate multiple VFMs [78].
We observe that Segment Anything 2 (SAM2) [48] offers an alternative paradigm. Designed for Video Object Segmentation, it operates as a prompt-and-propagate framework, where an object is specified via its mask and tracked across frames. To achieve this, SAM2 introduces a Memory Attention mechanism to implicitly match features across video frames and propagate masks over time with high spatial precision. While this feature matching is originally intended for object tracking based on visual similarity, we note that this architecture unifies two capabilities central to FSS within a single model: dense feature matching and high-quality mask generation. Building on this analogy, we propose repurposing the prompt-and-propagate framework to address the task of FSS, by reinterpreting the temporal dimension of videos as a collection of semantically related images.
@@
Addressing this inquiry means, in turn, to answer whether SAM2 has implicitly learned semantic-aware representations, despite its class agnostic pre-training. To answer this, we conduct a toy experiment in Fig. 2, testing SAM2 zero-shot performance on FSS datasets with different semantic variability. Interestingly, we observe that in low-semantic-shift scenarios, SAM2 achieves comparable or even higher performance with respect to state-of-the-art methods. However, on challenging datasets with high semantic shift, its performance drops drastically. A straightforward conclusion from these preliminary results would be that SAM2 may not have learned discriminative representations in its class-agnostic pretraining, making it unsuited for tasks requiring semantic alignment.
@@
We challenge this interpretation, based on the observation that SAM2 pretraining, focused on instance matching across frames, shares similarities with self-supervised learning frameworks [69, 68, 23, 5], which are known to elicit semantic understanding by enforcing feature invariance across views [9, 4]. Given this analogy, we posit that SAM2 does encode semantic information, which is however entangled with instance-specific features optimized for object tracking, reflecting experimental evidence in Fig. 2. If our intuition is true, it implies that i) this structure could be disentangled through lightweight transformations [1, 27, 32], such as adapter modules and ii) it should be learnable from a set of base classes and generalize to unseen categories [51, 13]. To this end, we i) intentionally use one of the simplest adapters in the literature, namely AdaptFormer [26], and ii) verify the generalization hypothesis through extensive experiments in strict few-shot benchmarks.
@@
Figure 1: We evaluate SAM2 in zero-shot on few-shot segmentation tasks on four datasets with varying degrees of semantic variability. On datasets [8, 14] with low semantic shift and high intra-class visual similarity, SAM2 matches or even outperforms state-of-the-art APSeg [22]. However, on more challenging datasets like COCO and LVIS, with high semantic shift (e.g., cruise ship vs. rowboat), its performance drops significantly, compared with GF-SAM [73].
@@
Figure 2: We extract SAM2 features from object instances across diverse images and visualize their distribution using the first three principal components from PCA. While zero-shot features from SAM2 lack clear semantic structure, after adapting features with SANSA, we observe the emergence of well-defined semantic clusters: semantically similar instances group together, forming coherent clusters despite intra-class variation in visual appearance.
@@
Building on this insight, we introduce SANSA (Semantically AligNed SegmentAnything 2), and show how to expose SAM2 latent semantic structure, repurposing its Memory Attention mechanism to shift from visual similarity to semantic similarity. The effectiveness of SANSA is illustrated in Fig. 2, where a 3D PCA, computed on unseen classes, reveals the emergent semantic organization of our features. We complement our method with a novel training objective designed to exploit SAM2 temporal continuity to convert each target image into a pseudo-annotation for subsequent frames.
@@
  • We are the first to investigate the semantic structure within SAM2 for the task of few-shot segmentation. We show that such semantics can be disentangled through bottleneck transformations, enabling a unified approach that reinterprets FSS as the task of tracking semantic concepts across images, by treating them as frames in a pseudo-video;
  • We validate SANSA through extensive experiments, achieving SOTA performance in strict FSS benchmarks while also outperforming generalist approaches in the ‘in-context’ scenario. Our experiments reveal that SAM2 encodes coarse-to-fine semantics, from high level concepts (e.g. dog vs. cat, +6.2% on COCO-20i) to fine-grained distinctions at category-level (e.g. dalmatian vs. bulldog, +9.3% on LVIS-92i) and part-level (e.g. hand vs. arm +5.1% on Pascal-Part);
@@ 2 Related works
Few-shot segmentation aims to segment a target image given an annotated reference. Early methods relied on compressed prototype representations [33, 62, 31, 36, 16], later replaced by attention- and correlation-based approaches [61, 74, 42, 41, 25] to better capture pixel-level relationships. More recently, research focused on leveraging the large-scale pretraining and generalization capabilities of foundation models [78, 77, 37, 64, 67]. Matcher [37] and GF-SAM [73] utilize a zero-shot pipeline: DINOv2 extracts correspondences which are used to prompt SAM for segmentation. Similarly, VRP-SAM [56] leverages a frozen SAM with an external encoder for feature matching. Recent works [77, 78, 39] explore the use of a single VFM [7] to jointly handle semantic understanding and mask prediction. DiffewS [78] exploits the emergent semantic correspondences in StableDiffusion [50] to unify the process, while SegIC [39] combines a frozen DINOv2 with a lightweight segmentation decoder. We posit that SAM2 implicitly encodes semantic information and has a natural propensity to generate accurate segmentation masks, and show that such latent semantic structure can be extracted and effectively repurposed to enable few-shot segmentation within a unified architecture.
@@
Semantic correspondences and Foundation Models. Finding correspondences between images is a longstanding problem in Computer Vision [6, 38, 28, 49, 44]. While early deep-learning approaches train dedicated models to establish semantic correspondences [28, 49], recent works [76, 57] have shown that Visual Foundation Models can provide robust semantic descriptors, enabling generalization across tasks [59, 2, 66]. Among them, DINOv2 [45] and StableDiffusion [50] have demonstrated a compelling ability to establish semantic correspondences between images [76, 44, 75]. Recently, Segment-Anything 2 [48] established itself as a foundation model for Video Object Segmentation. We observe that its pretraining, entailing matching object instances across frames under viewpoint changes and motion blur, closely parallels self-supervised learning frameworks [69, 68, 23] that derive semantic understanding through self-similarity training. However, the extent to which SAM2 embeddings encode (if any) semantic concepts has not been studied yet. Recent applications in specialized domains [80, 3] utilize a frozen SAM2 in low-semantic-shift scenarios (e.g. propagating masks across slices of 3D imagery given a support example). However, zero-shot SAM2 shows poor performance in standard few-shot segmentation benchmarks requiring higher level semantic understanding. In this work, we shed light on this matter, providing insights into SAM2 feature structure and showing that semantic content can be disentanged from its embeddings.
@@
Figure 3: Overview of SANSA: Given $k$ annotated reference images and a target image, we aim to produce the target mask by solving a semantic tracking problem. We construct a pseudo-video by concatenating reference and target images, then leverage SAM2 streaming pipeline to process reference frames and their annotations sequentially. We restructure SAM2 feature space to make its latent semantic structure explicit, enabling mask propagation based on semantic similarity from reference to target. The emergent semantic structure is visualized by the 3D PCA projection of $\mathcal{F}^{i}$.
@@ 3 Method
3.1 From Object Tracking to Semantic Tracking
[…]
  • Dense Feature Matching: Comprising the Memory Encoder, Memory Bank, and Memory Attention, this module establishes dense correspondences across frames. Concretely, given an object reference mask $\hat{y}_{r}$ (either predicted or given as prompt), the Memory Encoder constructs its memory representation, by fusing the mask with frame features $\mathcal{F}_{r}$:
    $\mathcal{I}_{r}=\mathcal{F}_{r}+\texttt{conv\_down}(\hat{y}_{r}).$ (1)
    This representation is stored in the Memory Bank. Subsequently, the features of target unannotated frames $\mathcal{F}_{t}$ undergo a cross-attention (Memory Attention) that aims at establishing dense correspondences between the current frame and the memory representations from previous frames:
    $\mathcal{F}^{i}_{t,\texttt{match}}=\text{Attention}\big{(}Q(\mathcal{F}_{t}^{i})K([\mathcal{I}_{r}^{0},...,\mathcal{I}_{r}^{i-1}])^{T}\big{)}V([\mathcal{I}_{r}^{0},...,\mathcal{I}_{r}^{i-1}]),$ (2)
    where $\mathcal{I}_{r}^{i}$ are past representations and $Q(\cdot),K(\cdot),V(\cdot)$ the query, key, and value projections.
  • High-quality Mask Generation: For unannotated frames, the Mask Decoder is tasked with refining the coarse features matches $\mathcal{F}^{i}_{t,\texttt{match}}$, which encode dense correspondences with prior object representations, to produce the segmentation output $\hat{y}_{i}$ (cf Fig. 3).
We observe that memory-based feature matching and mask decoding can be re-purposed for FSS, by reinterpreting the temporal dimension of video as a collection of semantically related images. In essence, rather than tracking a specific object across frames, we aim at tracking its semantic class.
Task Definition. We consider the general $k$-shot segmentation setting, where the model is given $K$ reference pairs $R={(x_{r}^{k},a_{r}^{k})}^{K}_{k=1}$, each consisting of an image $x_{r}^{k}\in\mathbb{R}^{H\times W\times 3}$ and its mask annotation $a_{r}^{k}\in[0,1]^{H\times W}$. Given a target image $x_{t}\in\mathbb{R}^{H\times W\times 3}$, the goal is to predict $y_{t}\in[0,1]^{H\times W}$, the segmentation mask of objects in $x_{t}$ that are semantically aligned with the reference. As shown in Fig. 3, we interpret the references and the target as a sequence of frames in a pseudo-video $\mathcal{M}$:
@@ 3.1 From Object Tracking to Semantic Tracking
[…]
3.2 Feature Adaptation
At the core of the feature matching mechanism lie the learned feature representations, central to the Memory Attention mechanism: the reference object representation $\mathcal{I}_{r}$ is constructed from $\mathcal{F}_{r}$ (Eq. 1), and matching is performed via cross-attention between target features $\mathcal{F}_{t}$ and $\mathcal{I}_{r}$ (Eq. 2).

### v2

@@ Abstract
Few-shot segmentation aims to segment unseen categories from just a handful of annotated examples. This requires mechanisms to identify semantically related objects across images and accurately produce masks. We note that Segment Anything 2 (SAM2), with its prompt-and-propagate mechanism, provides strong segmentation capabilities and a built-in feature matching process. However, we show that its representations are entangled with task-specific cues optimized for object tracking, which impairs its use for tasks requiring higher level semantic understanding. Our key insight is that, despite its class-agnostic pretraining, SAM2 already encodes rich semantic structure in its features. We propose SANSA (Semantically AligNed SegmentAnything 2), a framework that makes this latent structure explicit, and repurposes SAM2 for few-shot segmentation through minimal task-specific modifications. SANSA achieves state-of-the-art on few-shot segmentation benchmarks designed to assess generalization and outperforms generalist methods in the popular in-context setting. Additionally, it supports flexible promptable interaction via points, boxes, or scribbles, and remains significantly faster and more compact than prior approaches. Code at: https://github.com/ClaudiaCuttano/SANSA.
@@ 1 Introduction
Segmenting images is a core problem in Computer Vision, yet achieving high-quality results typically requires extensive human effort to annotate pixel-level masks. Moreover, conventional semantic segmentation methods  [chen2017deeplab, he2017mask, kirillov2018panoptic, cheng2022masked] struggle to generalize to unseen categories. Inspired by the human ability to recognize novel objects from just a few examples, few-shot segmentation (FSS) [li2021adaptive, wang2019panet, liu2022learning, shaban2017one] has emerged as a paradigm that leverages a small set of labeled reference samples to guide the segmentation of target images containing arbitrary, previously unseen classes.
To this end, recent work has turned to visual foundation models (VFMs) [bommasani2021opportunities], which offer rich visual representations and strong generalization capabilities [radford2021clip, yuan2023power]. A natural approach [liu2023matcher, zhang2024bridge, sun2024vrp] is to decouple the few-shot segmentation task into two stages: feature matching followed by promptable segmentation. This is typically achieved by combining DINOv2 [oquab2023dinov2], known for its strong semantic correspondence capabilities [zhang2023tale, meng2024segic, zhang2024telling, tang2023emergent], with Segment Anything [kirillov2023segment], which excels at producing high-quality segmentation masks [kirillov2023segment, zou2023segment, zhang2023personalize]. While effective, these modular approaches add computational overhead and require prompt engineering to coordinate multiple VFMs [zhu2024unleashing].
We observe that Segment Anything 2 (SAM2) [ravi2024sam] offers an alternative paradigm. Designed for Video Object Segmentation, it operates as a prompt-and-propagate framework, where an object is specified via its mask and tracked across frames. To achieve this, SAM2 introduces a Memory Attention mechanism to implicitly match features across video frames and propagate masks over time with high spatial precision. While this feature matching is originally intended for object tracking based on visual similarity, we note that this architecture inherently unifies two capabilities central to FSS within a single model: dense feature matching and high-quality mask generation. Building on this analogy, we propose repurposing the prompt-and-propagate framework to address the FSS task, by reinterpreting the temporal dimension of videos as a collection of semantically related images.
@@
Addressing this inquiry means, in turn, to answer whether SAM2 has implicitly learned semantic-aware representations, despite its class agnostic pre-training. To answer this, we conduct a toy experiment in Fig.˜2, testing SAM2 performance on FSS datasets with different semantic variability. Interestingly, we observe that in low-semantic-shift scenarios, SAM2 achieves comparable or even higher performance with respect to state-of-the-art methods. However, on challenging datasets with high semantic shift, its performance drops drastically. A straightforward conclusion from these preliminary results would be that SAM2 may not have learned discriminative representations in its class-agnostic pretraining, making it unsuited for tasks requiring semantic alignment.
@@
We challenge this interpretation, based on the observation that SAM2 pretraining, focused on instance matching across frames, shares similarities with self-supervised learning frameworks [yang2022inscon, yang2021instance, henaff2021efficient, bardes2022vicregl], which are known to elicit semantic understanding by enforcing feature invariance across views [caron2021emerging, bardes2021vicreg]. Given this analogy, we posit that SAM2 does encode semantic information, which is however entangled with instance-specific features optimized for object tracking, reflecting experimental evidence in Fig.˜2. If our intuition is true, it implies that i) this structure could be disentangled through lightweight transformations [aghajanyan2020intrinsic, hu2022lora, li2018measuring], such as adapter modules and ii) it should be learnable from a set of base classes and generalize to unseen categories [rusu2018meta, cheng2023meta]. To this end, we i) intentionally use one of the simplest adapters in the literature, namely AdaptFormer [houlsby2019adapter], and ii) verify the generalization hypothesis through extensive experiments in strict few-shot benchmarks.
@@
Figure 1: We evaluate frozen SAM2 on few-shot segmentation tasks on four datasets with varying degrees of semantic variability. On datasets [candemir2013lung, codella2019skin] with low semantic shift and high intra-class visual similarity, SAM2 matches or even outperforms state-of-the-art APSeg [he2024apseg]. However, on more challenging datasets like COCO and LVIS, with high semantic shift (e.g., cruise ship vs. rowboat), its performance drops significantly, compared with GF-SAM [zhang2024bridge]. The bottom row illustrates examples of ground-truth masks in both scenarios.
@@
Figure 2: We extract SAM2 features from object instances across diverse images and visualize their distribution using the first three principal components of PCA, mapped to RGB channels. The features appear entangled, with clusters mixing across categories, highlighting the lack of a coherent semantic structure in the original feature space. After adapting the feature space with SANSA, well-defined clusters emerge: semantically similar instances group together, forming coherent structures despite intra-class variation in visual appearance.
@@
Building on this insight, we introduce SANSA (Semantically AligNed SegmentAnything 2), and show how to expose SAM2 latent semantic structure, repurposing its Memory Attention mechanism to shift from visual similarity to semantic similarity. The effectiveness of SANSA is illustrated in Fig.˜2, where a 3D PCA, computed on unseen classes, reveals the emergent semantic organization of our features. We complement our method with a novel training objective designed to exploit SAM2 temporal continuity to convert each target image into a pseudo-annotation for subsequent frames.
@@
  • We are the first to investigate the semantic structure within SAM2. We show that such semantics can be disentangled through bottleneck transformations, enabling a unified approach that reinterprets few-shot segmentation as the task of tracking semantic concepts across images;
  • We validate SANSA through extensive experiments, achieving SOTA performance in strict FSS benchmarks while also outperforming generalist approaches in the ‘in-context’ scenario. Our experiments reveal that SAM2 encodes coarse-to-fine semantics, from high level concepts (e.g. dog vs. cat, +6.3% on COCO-20i) to fine-grained distinctions at category-level (e.g. dalmatian vs. bulldog, +8.3% on LVIS-92i) and part-level (e.g. hand vs. arm +4.6% on Pascal-Part);
@@ 2 Related works
Few-shot segmentation aims to segment a target image given an annotated reference. Early methods relied on compressed prototype representations [li2021adaptive, wang2019panet, lang2022learning, liu2022learning, fan2022self], later replaced by attention- and correlation-based approaches [wang2020few, zhang2021few, moon2023msi, Min_2021_ICCV, hong2022cost] to better capture pixel-level relationships. More recently, research focused on leveraging the large-scale pretraining and generalization capabilities of vision foundation models [zhu2024unleashing, zhang2023personalize, liu2023matcher, wang2023seggpt, xu2025hybrid]. Matcher [liu2023matcher] and GF-SAM [zhang2024bridge] utilize a training-free pipeline: DINOv2 extracts correspondences which are used to prompt SAM for segmentation. Similarly, VRP-SAM [sun2024vrp] leverages a frozen SAM with an external encoder for feature matching. Recent works [zhang2023personalize, zhu2024unleashing, meng2024segic] explore the use of a single VFM [bommasani2021opportunities] to jointly handle semantic understanding and mask prediction. DiffewS [zhu2024unleashing] exploit the emergent semantic correspondences in StableDiffusion [rombach2022diffusion] to unify the process, while SegIC [meng2024segic] combines a frozen DINOv2 with a lightweight segmentation decoder. Yet, recent findings [zhang2023tale, zhang2024telling] suggest that diffusion and DINOv2 features offer complementary but disjoint strengths: the former offers spatial precision but weak semantics, the latter strong semantics but sparse matches. In contrast, we posit that SAM2 unifies both properties: its features possess high spatial granularity and implicitly encode semantic information. We show that this latent semantic structure can be extracted, effectively enabling FSS within a unified architecture.
@@
Semantic correspondences and Foundation Models. Finding correspondences between images is a longstanding problem in Computer Vision [bay2006surf, lowe2004distinctive, kim2017fcss, rocco2018weakly, ofri2023neural]. While early deep-learning approaches train dedicated models to establish semantic correspondences [kim2017fcss, rocco2018weakly], recent works [zhang2023tale, tang2023emergent] have shown that VFMs enable generalization across tasks [tumanyan2022splicing, awais2025foundation, wei2024stronger, cuttano2025samwise]. Among them, DINOv2 [oquab2023dinov2] and StableDiffusion [rombach2022diffusion] have demonstrated a compelling ability to establish semantic correspondences between images [zhang2023tale, ofri2023neural, zhang2024telling]. Recently, Segment Anything 2 [ravi2024sam] established itself as a foundation model for Video Object Segmentation. We observe that its pretraining, entailing matching object instances across frames under viewpoint changes and motion blur, closely parallels self-supervised learning frameworks [yang2022inscon, yang2021instance, henaff2021efficient] that derive semantic understanding through self-similarity training. However, the extent to which SAM2 embeddings encode (if any) semantic concepts has not been studied yet. Recent applications in specialized domains [zu2025medicalsam, bai2024medicalsam] utilize a frozen SAM2 in low-semantic-shift scenarios (e.g. propagating masks across slices of 3D imagery given a support example). However, frozen SAM2 shows poor performance in standard FSS benchmarks requiring higher level semantic understanding. In this work, we shed light on this matter, providing insights into SAM2 feature structure and showing that semantic content can be disentanged from its embeddings.
@@
Figure 3: Overview of SANSA: Given $k$ annotated reference images and a target image, we construct a pseudo-video by concatenating them, then leverage SAM2 streaming pipeline to process reference frames together with their annotations sequentially. We restructure SAM2 feature space to make its latent semantic structure explicit, enabling mask propagation based on semantic similarity from reference to target. The emergent semantic structure is visualized by the 3D PCA projection of $\mathcal{F}$.
@@ 3 Method
Task Definition. We consider the general $k$-shot segmentation setting, where the model is given $K$ reference pairs $R={(x_{r}^{k},a_{r}^{k})}^{K}_{k=1}$, each consisting of an image $x_{r}^{k}\in\mathbb{R}^{H\times W\times 3}$ and its mask annotation $a_{r}^{k}\in[0,1]^{H\times W}$. Given a target image $x_{t}\in\mathbb{R}^{H\times W\times 3}$, the goal is to predict $y_{t}\in[0,1]^{H\times W}$, the segmentation mask of objects in $x_{t}$ that are semantically aligned with the reference. As shown in Fig.˜3, we interpret the references and the target as a sequence of frames in a pseudo-video $\mathcal{M}$:
@@
3.1 From Object Tracking to Semantic Tracking with SAM2
Segment Anything Model 2 extends SAM [kirillov2023segment] to Promptable Video Object Segmentation. Like its predecessor, SAM2 comprises three main elements: an Image Encoder, a Prompt Encoder, and a Mask Decoder. Specifically, given an image $x_{r}^{k}$ with features $\mathcal{F}_{r}^{k}$, and a prompt $a_{r}^{k}\in\big\{\texttt{mask, point, box}\big\}$, the Mask Decoder processes them to produce the segmentation mask $\hat{y}_{r}^{k}$. The key innovation of SAM2 lies in its extension to video domain: masks can be propagated across new unannotated frames $x_{t}$ without additional prompts, thanks to a memory mechanism. As shown in Fig.˜3, we conceptually decompose SAM2 architecture into two functional components:
  • Dense Feature Matching: Comprising the Memory Encoder, Memory Bank, and Memory Attention, this module establishes dense correspondences across frames. Concretely, given an object reference mask $\hat{y}_{r}^{k}$ (either predicted or given as prompt), the Memory Encoder constructs its memory representation, by fusing the mask with frame features $\mathcal{F}_{r}^{k}$:
    $\mathcal{I}_{r}^{k}=\mathcal{F}_{r}^{k}+\texttt{conv\_down}(\hat{y}_{r}^{k}).$ (2)
    This representation is stored in the Memory Bank. Subsequently, the features $\mathcal{F}_{t}$ of target unannotated frames undergo a cross-attention (Memory Attention) that aims at establishing dense correspondences between the current frame and the memory representations from previous frames:
    $\mathcal{F}_{t,\texttt{match}}=\text{Attention}\big(Q(\mathcal{F}_{t})K([\mathcal{I}_{r}^{0},...,\mathcal{I}_{r}^{k}])^{T}\big)V([\mathcal{I}_{r}^{0},...,\mathcal{I}_{r}^{k}]),$ (3)
    where $\mathcal{I}_{r}^{k}$ are past representations and $Q(\cdot),K(\cdot),V(\cdot)$ the query, key, and value projections.
  • High-quality Mask Generation: For unannotated frames, the Mask Decoder is tasked with refining the coarse features matches $\mathcal{F}_{t,\texttt{match}}$, which encode dense correspondences with prior object representations, to produce the segmentation output $\hat{y}_{t}$ (cf Fig.˜3).
We propose to repurpose the memory-based feature matching and mask decoding mechanisms, reinterpreting the temporal dimension of videos as a collection of semantically related images. Thus, rather than tracking a specific object across continuous frames, we aim at tracking its semantic class.
We highlight two advantages of this formulation: first, unlike recent approaches [sun2024vrp, meng2024segic, zhu2024unleashing], our model naturally supports variable $k$-shots without modifications; second, our solution seamlessly supports promptable FSS, where prompts can take the form $a_{r}^{k}\in\big\{\texttt{mask, point, box, scribble}\big\}$, removing the reliance on pixel-level annotations. Finally, we note that reference frames are encoded in Memory Bank without undergoing Memory Attention (cf Fig.˜3). By avoiding cross-referencing, we ensure predictions for the target image to be invariant to the ordering of reference images.
3.2 SAM2 Feature Adaptation
At the core of the feature matching mechanism lie the learned feature representations, central to the Memory Attention mechanism: the reference object representation $\mathcal{I}_{r}$ is constructed from $\mathcal{F}_{r}$ (Eq.˜2), and matching is performed via cross-attention between target features $\mathcal{F}_{t}$ and $\mathcal{I}_{r}$ (Eq.˜3).